OpenStudy (anonymous):

Can someone help me with this L'Hopital problem: lim as x approaches 1 from the right or (lnx)^(x-1)

OpenStudy (anonymous):

I was thinking you could do y=lim.... and then take ln of both sides. But that just makes it more difficult.

OpenStudy (anonymous):

take the log first then do it

OpenStudy (anonymous):

yes u have to remove the (x-1) from the exponent first

OpenStudy (anonymous):

or else rewrite as \[\large \ln(x)^{x-1}=e^{(x-1)\ln(\ln(x))}\]

OpenStudy (anonymous):

either way the work is identical, finding \[\lim_{x\to 1^+}(x-1)\ln(\ln(x))\]

OpenStudy (amistre64):

\[[ln(x)]^{x-1}=ln^x(x)*ln(x)\]

OpenStudy (amistre64):

divide, not multiply

OpenStudy (amistre64):

\[[ln(x)]^{x-1}=ln^x(x)~/~ln(x)\]

OpenStudy (amistre64):

right?

OpenStudy (anonymous):

yeah, but my guess is that it is now going to be even more complicated although i could be way off

OpenStudy (amistre64):

was this your suggestion? \[y=(lnx)^{(x-1)}\]\[e^y=e^{(x-1)~lnx}\]

OpenStudy (anonymous):

maybe not though...

OpenStudy (anonymous):

no it was just two rewrite \(b^x\) as \(e^{x\ln(b)}\)

OpenStudy (anonymous):

usually math teacher says "take the log" and "simplify using properties of the log" then compute the limit and then since you take the log as a first step you "exponentiate" (if that is a word) at the end

OpenStudy (anonymous):

so in this problem it would be \[y=\ln(x)^{x-1}\] \[\ln(y)=\ln(\ln(x)^{x-1})=(x-1)\ln(x)\]

OpenStudy (anonymous):

oops i mean \[(x-1)\ln(\ln(x))\]

OpenStudy (amistre64):

my precalc teacher told us that whenever you see a convoluted setup, the answer is either going to be 0 or 1 :)

OpenStudy (anonymous):

lol what about \[\lim_{x\to \infty}(1+\frac{3}{x})^x\]?

OpenStudy (anonymous):

but back to my screed, this is all math teacher nonsense

OpenStudy (anonymous):

since by definition \[b^x:=e^{x\ln(b)}\]

OpenStudy (anonymous):

and so by definition \[\large \ln(x)^{x-1}=e^{(x-1)\ln(\ln(x))}\]

OpenStudy (anonymous):

all the work is still computing the limit in the sky however, so it really makes no difference

OpenStudy (amistre64):

i wonder about a power series, but that thought really hasnt formed and im sure it would be a nightmare

OpenStudy (anonymous):

eeeew

OpenStudy (anonymous):

seems nightmareish

OpenStudy (amistre64):

\[ln(x)=\sum_1\frac{-(1-x)^n}{n}\] \[[ln(x)]^{x-1}=\sum_1\left[\frac{-(1-x)^n}{n}\right]^{x-1}\]

OpenStudy (anonymous):

yikes!!

OpenStudy (amistre64):

oh it aint that bad :) at x=1 that zeros out to 1 i believe

OpenStudy (anonymous):

My book kinda of gives me an answer, it say to do: Let y= lim ln(x)^(x-1), then somehow its goes to = lim (x-1)ln(x) =0 Hence the lim is equal to 1

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

it doesn't go to \((x-1)\ln(\ln(x))\) that is what you get when you take the log

OpenStudy (anonymous):

no, its doesnt even say to take ln of both sides

OpenStudy (anonymous):

which is unsusal, beacuse the soulutions to the other problems like this one actually say take ln of both sides

OpenStudy (anonymous):

ok then maybe i misread the problem could it be \[\large \ln(x^{x-1})\]

OpenStudy (anonymous):

which is not the same as \[\large (\ln(x))^{x-1}\] at all

OpenStudy (anonymous):

nope, it's